# Modeling Future Conversation Turns to Teach LLMs to Ask Clarifying Questions

### v1

@@ 5 Results
Table 1: Main results. Our method (Clarify DPO (Ours)) yields the optimal results in Answer F1 for all base models at the cost of always asking clarifying questions. Our method that selectively asks clarifying questions (Clarify-or-Direct-Ans DPO (Ours)) strikes a balance between efficiency (Avg Turns) and effectiveness (Answer F1). Bold results denote statistically significant differences from all other methods with the same base model (over $N=10,000$ samples with $p<0.01$).
Base Turn 2 Model Inference Avg Turns ($\downarrow$) Answer F1 ($\uparrow$)
Unambig Ambig Total
Llama2 (7b) Direct-Ans SFT Greedy 1 0.254 0.168 0.211
Direct-Ans SFT Sampled 1 0.250 0.172 0.214
Clarify SFT Greedy 2 0.310 0.216 0.259
Clarify DPO (Starling) Greedy 2 0.310 0.257 0.283
Clarify DPO (Ours) Greedy 2 0.383 0.282 0.328
Clarify-or-Direct-Ans DPO (Ours) Greedy 1.56 0.289 0.211 0.243
Gemma (7b) Direct-Ans SFT Greedy 1 0.254 0.168 0.211
Direct-Ans SFT Sampled 1 0.250 0.172 0.214
Clarify SFT Greedy 2 0.357 0.236 0.288
Clarify DPO (Starling) Greedy 2 0.339 0.257 0.295
Clarify DPO (Ours) Greedy 2 0.407 0.286 0.339
Clarify-or-Direct-Ans DPO (Ours) Greedy 1.61 0.282 0.222 0.246
Llama3 (8b) Direct-Ans Greedy 1 0.312 0.192 0.248
Direct-Ans Sampled 1 0.282 0.202 0.247
Clarify SFT Greedy 2 0.376 0.265 0.315
Clarify DPO (Starling) Greedy 2 0.362 0.267 0.309
Clarify DPO (Ours) Greedy 2 0.429 0.317 0.365
Clarify-or-Direct-Ans DPO (Ours) Greedy 1.57 0.352 0.251 0.291
@@
We report our main results in Table 1. Overall, systems that are equipped to ask clarifying questions outperform Direct-Ans models in effectiveness (Answer F1). Clarify-or-Answer methods strike a balance between effectiveness and efficiency. We further investigate the behaviors of Clarify-or-Answer methods in Section 5.1.
@@
We also find that training with double-turn preference labeling scheme consistently yields the best answer F1 overall. We observe mixed results from training with single-turn RLHF annotations from Starling: yielding positive results with Llama2 as the base LLM and minor increases/decreases in performance for Gemma and Llama3 when evaluated over both ambiguous and unambiguous queries. These results support our claim that single-turn RLHF annotation struggles to identify useful clarifying questions. In contrast, training with double-turn preferences, where clarifying questions are assessed based on their later outcomes, demonstrates consistent improvements in the LLM’s ability to generate useful clarifying questions, on both ambiguous and unambiguous queries. This demonstrates that clarifying questions not only help models disambiguate user intents for ambiguous queries, but also can help models recover correct answers in general even for unambiguous queries. We include examples of generated clarifying questions in Appendix C.
@@
Table 2: Ablations comparing training Clarify DPO (Ours) systems using SFT training on clarifying questions generated from $Y_{\text{human}}$ or $Y_{\text{model}}$ answer sets or both.
Base Model SFT Dataset Unambig Ambig Full
Llama2 (7b) Human 0.288 0.236 0.259
Model 0.299 0.264 0.283
Both 0.383 0.282 0.328
Gemma (7b) Human 0.379 0.274 0.323
Model 0.391 0.283 0.329
Both 0.407 0.286 0.339
Llama3 (8b) Human 0.401 0.309 0.349
Model 0.423 0.312 0.364
Both 0.429 0.317 0.365
@@
Ablations: Comparing Clairifying Question Generated with Human Annotated Answers vs. Model Predicted Answers
@@
In our main experiments, we performed SFT training on the union of clarifying questions generated from model-identified $(Y_{\text{model}})$ and human-identified $(Y_{\text{human}})$ feasible answer sets. Is one of them more useful than the other? In Table 2, we compare performance using SFT examples constructed from our $Y_{\text{model}}$ and $Y_{\text{human}}$ answer sets, after RLHF training with double-turn preference. Looking first at our results using Llama2, we see clear improvements when using our model-ambiguity SFT dataset over using our human-ambiguity dataset alone. Surprisingly, we find that these improvements even hold when evaluated against the answer sets from AmbigQA, despite the fact that our human-ambiguity SFT dataset was generated to distinguish between answers labeled in the AmbigQA training split.
@@
Next, looking at our results using Gemma and LLama3, we find that there are only minor overall improvements from training when our model-ambiguity SFT dataset. Recall that our model-ambiguity dataset was using feasible answer sets $Y_{\text{model}}$ from few-shot Llama2 predictions. These results, therefore, suggest that constructing a model-ambiguity dataset that is specific to the base model improves the resultant clarifying question quality.
@@ A Joint Model for Asking Clarifying Questions and Predicting Answers
So far, our full systems have two separate LLMs. Here, we experiment with using a single LLM for both the clarification $(x\rightarrow q)$ and answer $(x,q,a_{i}\rightarrow y_{i})$ conversation turns. To create this joint model, we take inspiration from prior works demonstrating that averaging the weights of multiple finetuned models is a simple and effective approach to joining the capabilities of finetuned models into a single LLM (Wortsman et al., 2022). We merge the LoRA parameter updates from our clarification model (trained with preferences from full interactions) and the answer prediction models used in our main results (Table 1).
Table 3: Results comparing F1 using separate clarifying question and answering models and using a single joint model to perform both tasks.
Base LLM Cl. & Ans. Models Unambig Ambig Full
Llama2 (7b) Separate 0.383 0.282 0.328
Joint 0.324 0.250 0.287
Gemma (7b) Separate 0.407 0.286 0.339
Joint 0.362 0.274 0.314
Llama3 (8b) Separate 0.429 0.317 0.365
Joint 0.410 0.292 0.346
We report our results from these experiments in Table 3, where we compare performance against using separate LLMs for each conversation turn. We see that overall performance degrades when using a joint model for both turns; however, we are still able to retain most of the performance gains over Direct-Ans and QA-with-Clarification methods trained with single-turn preferences. Future work may explore alternative training methods that mitigate the performance degradation, such as using mixture-of-experts methods for LLMs (Jiang et al., 2024) or improved multi-turn learning algorithms (Zhou et al., 2024b).
@@ 5.1 Determining When Clarifying Questions are Necessary
In this section, we further investigate our system’s ability to identify when clarifying questions are necessary. To evaluate our Clarify-or-Answer model’s predictions, we report Direct-Answer Accuracy: Did the the system correctly decide to directly answer the question when the question was unambiguous and the greedy predicted direct-answer was correct. To determine whether the greedy predicted direct-answer was correct, we decode our Clarify-or-Answer system’s greedy direct-answer prediction for each input question.
We compare our models predictions (C-or-A-Pred) against the random baseline for determining whether to directly-answer or ask a clarifying question. To compute this, we fix the percent of direct-answer responses (DA%) to the model’s clarify-or-answer predictions and randomly sample DAof predictions to directly answer, and predict clarifying questions for the remaining input queries.
Table 4: Results for Clarify-or-Answer methods. For each model, we include the percent of questions directly answered (DA%), and evaluate using F1, and Direct-Answer Acc.
Model Method F1 DA Acc
Llama2 (DA 44%) Random 0.234 0.554
C-or-A Pred 0.243 0.619
Gemma (DA 38%) Random 0.237 0.604
C-or-A Pred 0.246 0.643
Llama3 (DA 43%) Random 0.280 0.549
C-or-A Pred 0.365 0.595
Results
We report our results from our Clarify-or-Answer systems in Table 4. Our Clarify-or-Answer system’s predictions, however, are able to consistently outperform random baseline, with the biggest improvement being from the Direct-Answer accuracy metric. Prior work has demonstrated that the similar task of classifying ambiguous from unambiguous questions is a challenging task, with the best LLM-based methods achieving 58% accuracy (Cole et al., 2023). While our task is slightly different, determining if the question is ambiguous or the direct-answer response is incorrect, we find that this task is similarly difficult for LLM systems. The relatively large gains on Direct-Answer accuracy compared to the F1, indicates that we can expect performance on end-task metrics to further improve with better clarifying question generations.
@@ 6 Related Work
Clarifying Question Generation Prior works for training LLMs to ask clarifying questions have focused on settings where the types of clarification required are fixed by the task (i.e., there is fixed set of relevant input features (Wei et al., 2018; Kuhn et al., 2022)) or the input (e.g., reading-comprehension tasks where the context contains multiple correct answers (Guo et al., 2021)). Such works have studied clarifying question generation for a variety of tasks, including gauging social and moral situations (Pyatkin et al., 2022) and image classification (Yu et al., 2019). Andukuri et al. (2024) trained LLMs to ask clarifying questions to resolve task ambiguity, where a single request may have multiple different indented tasks depending on the user. In this work, the authors simulate ideal responses from a fixed pool of simulated personas, and task models with asking clarifying questions to generate a tailored response. Hong et al. (2023) also uses LLM-simulated dialogues for training. Recent work has also noted the scarcity of clarifying questions in responses from existing LLMs, and have studied promoting methods for eliciting such responses with greater frequency (Deng et al., 2023; Shaikh et al., 2023). Li et al. (2023) found that clarifying questions can be useful for learning individualized preferences.
@@ 7 Limitations and Future Work
Throughout this work, we only consider systems for one or two turn interactions and does not consider cases where asking additional clarifying questions might be helpful after observing the user’s clarifying answer. Future work might explore methods for extending our evaluation framework and double-turn preference annotation methods to accommodate general multi-turn interactions with users. Furthermore, the systems explored in this work also do not model dialogue acts (Stolcke et al., 2000) outside of predicting a single answer or asking a clarifying question. At times, it might be more appropriate for LLMs to generate responses for different behaviors (e.g., Overton responses containing multiple answers if there are only a few possible interpretations. Long vs. short answers for complex queries. Abstaining in lieu of predicting erroneous answers). Future work might consider how clarifying questions should be used in conjunction with other strategies.
Finally, throughout this work we rely on simulated user interactions for labeling double-turn preferences and for evaluation. While future work might consider applying our evaluation framework and double-turn preference annotation scheme to human annotators, there are several challenging hurdles. In particular, annotating a clarifying answer to an input query, clarifying question, and known expected answer can be challenging, particularly in open-domain QA where queries often ask about niche topics. Furthermore, relying on annotators to simply provide their best own answers to clarifying questions may lead issues in coverage of possible query interpretations. These challenges indicate that relying on human evaluators for clarifying questions can be costly and its reliability is highly dependent on the pool of annotators.
@@ 8 Conclusion
We propose a method for training LLMs to ask clarifying questions with double-turn preferences. To evaluate our systems, we develop an automatic framework for evaluating systems that ask clarifying questions using user simulated user interactions. Our results on Open-Domain QA demonstrate that training systems to ask clarifying questions with double-turn preferences improves QA performance over training with standard single-turn preference labels. We also demonstrate that training with double-turn preferences be used to train models to judiciously decide when to ask for clarification.
@@ Acknowledgments
We thank the UT NLP community for providing feedback throughout the project. This work was in part supported by Cisco Research. Any opinions, findings and conclusions, or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of Cisco Research. The work is partially supported by a gift from Google Research. This work has taken place in part in the Rewarding Lab at UT Austin. The Rewarding Lab is supported by NSF (IIS-2402650), ONR (N00014-22-1-2204), EA Ventures, Bosch, and UT Austin’s Good Systems grand challenge.

### v2

@@ 5 Results
Table 1: Main results, separated by systems that never clarify (# Turns = 1), always clarify (# Turns = 2), or can do both (# Turns $\in[1,2]$). The shaded blocks represent systems that simulate future turns. Our method (Clarify DPO w/ Match) yields the optimal results in Answer F1 at the cost of always asking clarifying questions. Our method that selectively asks clarifying questions (Clarify-or-Direct-Ans DPO strikes a balance between efficiency (# Turns) and effectiveness (Answer F1). Bold results denote statistically significant differences from all other methods with the same base model (over $N=10,000$ samples with $p<0.01$).
Llama2 Llama3 Gemma
# Answer F1 ($\uparrow$) # Answer F1 ($\uparrow$) # Answer F1 ($\uparrow$)
($\downarrow$)↓ ) Unamb / Amb / All ($\downarrow$) Unamb / Amb / All ($\downarrow$) Unamb / Amb / All
Direct-Ans SFT
   w/ Greedy 1 25.4 / 16.8 / 21.1 1 31.2 / 19.2 / 24.8 1 26.1 / 16.8 / 21.1
   w/ Sampled 1 25.0 / 17.2 / 21.4 1 28.2 / 20.2 / 24.7 1 23.7 / 17.9 / 21.4
Clarify SFT 2 31.0 / 21.6 / 25.9 2 37.6 / 26.5 / 31.5 2 35.7 / 23.6 / 28.8
Clarify DPO
   w/ RM 2 31.0 / 25.7 / 28.3 2 36.2 / 26.7 / 30.9 2 33.9 / 25.7 / 29.5
   w/ Likelihood 2 30.2 / 23.9 / 27.2 2 43.5 / 29.6 / 359 2 37.3 / 26.8 / 31.5
   w/ Match 2 38.3 / 28.2 / 32.8 2 42.9 / 3.17 / 36.5 2 40.7 / 28.6 / 33.9
Clarify-or-Direct-Ans
   SFT 1.12 25.6 / 18.4 / 21.3 1.40 35.3 / 23.5 / 28.2 1.43 22.3 / 19.0 / 20.3
   DPO 1.56 28.9 / 21.1 / 24.3 1.57 35.2 / 25.1 / 29.1 1.61 28.2 / 22.2 / 24.6
@@
We report our main results in Table 1, capturing the number of user turns in the conversation (# $\downarrow$) and Answer F1. Overall, systems that are equipped to ask clarifying questions outperform Direct-Ans models in effectiveness (Answer F1). Clarify-or-Answer methods strike a balance between effectiveness and efficiency.
@@
Training with our double-turn preference labeling scheme consistently achives the best answer F1, with our match method yielding the best results overall. We observe mixed results from training with single-turn RLHF annotations from the Starling RM: yielding positive results with Llama2 as the base LLM and minor increases/decreases in performance for Gemma and Llama3 when evaluated over both ambiguous and unambiguous queries. These results support our claim that single-turn RLHF annotation struggles to identify useful clarifying questions. In contrast, training with double-turn preferences, where clarifying questions are assessed based on their later outcomes, demonstrates consistent improvements in the LLM’s ability to generate useful clarifying questions, on both ambiguous and unambiguous queries. This demonstrates that clarifying questions not only help models disambiguate user intents for ambiguous queries, but also can help models recover correct answers in general even for unambiguous queries. We include such example generations in Appendix D.
@@
Table 2: Answer F1 results comparing training Clarify DPO (Ours) systems using SFT training on clarifying questions generated from $Y_{\text{human}}$ or $Y_{\text{model}}$ answer sets or both.
SFT Data Llama2 (7b) Llama3 (7b) Gemma (7b)
Unamb Amb All Unamb Amb All Unamb Amb All
Human 28.8 23.6 25.9 40.1 30.9 34.9 37.9 27.4 32.3
Model 29.9 26.4 28.3 42.3 31.2 36.4 39.1 28.3 32.9
Both 38.3 28.2 32.8 42.9 31.7 36.5 40.7 28.6 33.9
@@
Ablations: Comparing Clarifying Question Generated with Human Annotated Answers vs. Model Predicted Answers
@@
So far, we performed SFT training on the union of clarifying questions generated from model-identified $(Y_{\text{model}})$ and human-identified $(Y_{\text{human}})$ feasible answer sets. Is one of them more useful than the other? In Table 2, we compare performance using SFT examples constructed from our $Y_{\text{model}}$ and $Y_{\text{human}}$ answer sets, after RLHF training with double-turn preference.
@@
With Llama2 base model, we see clear improvements when using our model-ambiguity SFT dataset over using our human-ambiguity dataset alone. These improvements even hold when evaluated against the answer sets from AmbigQA, despite the fact that our human-ambiguity SFT dataset was generated to distinguish between answers labeled in the AmbigQA training split. However, on Gemma and LLama3, we find only minor improvements from training with model-ambiguity SFT dataset. The model-ambiguity dataset was using feasible answer sets $Y_{\text{model}}$ from Llama2 predictions. This gap suggests that constructing a model-ambiguity dataset that is specific to the base model improves the resultant clarifying question quality.
@@ A Joint Model for Asking Clarifying Questions and Predicting Answers
Here,
Table 3: Answer F1 results comparing separate versus joint clarifying question and answering models (Cl / Ans).
Llama2 (7b) Llama3 (7b) Gemma (7b)
Unamb Amb All Unamb Amb All Unamb Amb All
Separate 38.3 28.2 32.8 42.9 31.7 36.5 40.7 28.6 33.9
Joint 32.4 25.0 28.7 41.0 29.2 34.6 36.2 27.4 31.4
we experiment with using a single LLM for both the clarification $(x\rightarrow q)$ and answer $(x,q,a_{i}\rightarrow y_{i})$ conversation turns. To create this joint model, we take inspiration from prior works demonstrating that averaging the weights of multiple finetuned models is a simple and effective approach to joining the capabilities of finetuned models into a single LLM (Wortsman et al., 2022). We merge the LoRA parameter updates from our clarification model (trained with preferences from full interactions) and the answer prediction models used in our main results (Table 1).
Table 3 compares the joint model against using separate LLMs for each conversation turn. Overall performance degrades when using a joint model for both turns; however, we are still able to retain most of the performance gains over Direct-Ans and QA-with-Clarification methods trained with single-turn preferences. Future work may explore alternative training methods that mitigate the performance degradation, such as using mixture-of-experts methods for LLMs (Jiang et al., 2024) or improved multi-turn learning algorithms (Zhou et al., 2024b).
@@ 5.1 Determining When Clarifying Questions are Necessary
We further investigate our system’s ability to identify when clarifying questions are necessary. To evaluate our Clarify-or-Answer model’s predictions, we report Direct-Answer Accuracy: Did the the system correctly decide to directly answer the question when the question was unambiguous and the greedy predicted direct-answer was correct. To determine whether the greedy predicted direct-answer was correct, we decode our Clarify-or-Answer system’s greedy direct-answer prediction for each input question. We also report Ambig Acc: the accuracy of a system’s clarify-vs-answer predictions on the human ambiguous-vs-unambiguous labels.
Table 4: Additional results for Clarify-or-Answer method (Llama2 model). We include the percent of questions directly answered (DA%), and evaluate using F1, Direct-Answer (DA) Acc., and Ambig (A) Acc.
Method (DA%) Ans F1 $\uparrow$ DA Acc $\uparrow$ A Acc $\uparrow$
Random (44%) 23.4 55.4 52.1
Ours (44%) 24.3 61.9 53.7
Ours w/ ProCoT (35%) 15.6 60.1 49.2
Ours w/ PPDPP (46%) 23.9 57.7 59.0
We compare our models predictions (C-or-A-Pred) against a random baseline and two additional methods, ProCot and PPDPP. To compute the random baseline, we fix the percent of direct-answer responses (DA%) to our model’s clarify-or-answer predictions and randomly sample DAof predictions to directly answer, and predict clarifying questions for the remaining queries. ProCot Deng et al. (2023a) utilizes a CoT prompt to determine whether or not the input is ambiguous and a clarifying question is necessary. In practice, we use the instruction-tuned variants of each base model to make these predictions, as we find irregular behavior when applying the CoT prompt to non-instruction tuned models. PPDPP Deng et al. (2023b) proposes a method for training an auxiliary model to predict how models should respond. Here, we train an additional copy of the base model make the binary prediction of whether asking a clarifying question is necessary for the system recover more all answers. ProCot and PPDPP do not train models to what questions to ask, and are only designed to predict whether clarifying is necessary. To evaluate these baselines, we pass their predictions to our trained Clarify-or-Answer model, which then generate the clarifying question of direct answer according to the baseline’s predicitons. We include implementation details in Appendix B.
We report our results with Llama2 in Table 4 (results with other base models in Appendix B). Our Clarify-or-Answer system’s predictions consistently outperforms all baselines across all metrics except Ambig Accuracy. Despite using a stronger base model (instruction tuned variant instead of base model), the prompt-based method (ProCot Prompt) performs the worst, as it is not trained in-domain. Training a separate model for ambiguity prediction (PPDPP) performs well, showing almost 10% higher ambiguity accuracy than our full model. However, our method still outperforms this model in answer accuracy. This demonstrates that our method (1) more frequently asks clarifying questions when it’s useful to do so; and (2) avoids asking for clarification on ambiguous questions when the model is not able to successfully use the clarifying question. The relatively large gains on Direct-Answer accuracy compared to F1 indicate that we can expect performance on end-task metrics to further improve with better clarifying question generation.
@@ 6 Related Work
Clarifying Question Prior works for training LLMs to ask clarifying questions have focused on settings where the types of clarification required are fixed by the task (i.e., there is fixed set of relevant input features (Wei et al., 2018; Kuhn et al., 2022)) or the input (e.g., reading-comprehension tasks where the context contains multiple correct answers (Guo et al., 2021)). Such works have studied clarifying question generation for a variety of tasks, including gauging social and moral situations (Pyatkin et al., 2022) and image classification (Yu et al., 2019).
Closely related to our work, Andukuri et al. (2024) trained LLMs to ask clarifying questions to resolve task ambiguity, where a single request may have multiple different indented tasks depending on the user, after user simulation. Compared to their experimental setting of simulated user persona for open-ended tasks, we address real-world open domain question answering task. Furthermore, we leverage preference optimization framework while they focus on generating supervised fine-tuning data by filtering low-quality clarifying questions. Hong et al. (2023) also uses LLM-simulated dialogues for training. Recent work has also noted the scarcity of clarifying questions in responses from existing LLMs, and have studied promoting methods for eliciting such responses with greater frequency (Deng et al., 2023a; Shaikh et al., 2023). Li et al. (2023) found that clarifying questions can be useful for learning individualized preferences.
@@ 7 Limitations and Future Work
We only consider systems for one or two turn interactions and does not consider cases where asking additional clarifying questions might be helpful after observing the user’s clarifying answer. Future work might explore methods for extending our evaluation framework and double-turn preference annotation methods to accommodate general multi-turn interactions with users. Furthermore, the systems explored in this work also do not model dialogue acts (Stolcke et al., 2000) outside of predicting a single answer or asking a clarifying question. At times, it might be more appropriate for LLMs to generate responses for different behaviors (e.g., Overton responses containing multiple answers if there are only a few possible interpretations. Long vs. short answers for complex queries. Abstaining in lieu of predicting erroneous answers). Future work might consider how clarifying questions should be used in conjunction with other strategies. Another important direction is extending double-turn preference method for other tasks. Our evaluation framework is intended to be easily adaptable, only requiring (1) multiple annotator labels for expected outputs and (2) an changing end-task evaluation metric for the new setting.
@@ 8 Conclusion
We propose a method for training LLMs to ask clarifying questions with double-turn preferences and an an automatic framework for evaluating systems that ask clarifying questions using user simulated user interactions. Our QA results demonstrate that training systems to ask clarifying questions with double-turn preferences improves performance over training with standard single-turn preference labels and can be used to train models to judiciously decide when to ask for clarification.
@@ Acknowledgments
We thank the UT NLP community for providing feedback throughout the project. This work was in part supported by Cisco Research. Any opinions, findings and conclusions, or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of Cisco Research. The work is partially supported by a gift from Google Research. This work has taken place in part in the Rewarding Lab at UT Austin. The Rewarding Lab is supported by NSF (IIS-2402650), ONR (N00014-22-1-2204), EA Ventures, Bosch, UT Austin’s Good Systems grand challenge, and Open Philanthropy. This work was done in part while the first and last author was visiting the Simons Institute for the Theory of Computing.
@@ Appendix A Dataset Details
A.1 Human and Model Ambiguity SFT and RLHF Training Dataset Details
Appendix B Additional Experiments
B.1 Clarify-or-Answer Baselines
Here, we describe the additional baselines introduced in Section 5.1.
Proactive chain-of-thought (ProCoT) prompt Deng et al. (2023a) presents a chain-of-thought prompt method for instructing LLMs to first determine whether or not an input question is ambiguous before generating a direct-answer or clarifying question response. The need for instruction-following abilities for this baseline necessitates using the instruction-tuned variants of each base model. Therefore, performance of this approach may be inflated due to the stronger base model.
Plug-and-Play Dialogue Policy Planner (PPDPP) Deng et al. (2023b) proposes a SFT and RL methods for training a separate model predicting the proper dialogue act for the next assistant turn. Because their method is not designed to train systems to generate responses, only to decide which type or response to predict, we train a separate PPDPP model from our base LLM via SFT. This system is trained over the SFT training sets over the gold dialogue act: clarify if the question is ambiguous or directly answer otherwise. We do not include their RL method, as the authors report comparable performance with and without additional RL training. We found training to be hyper parameter sensitive, with the suggested hyperparameter training systems to always select a single option. Using this trained system, we then use its predictions to determine whether to predict a clarifying question or direct-answer response using our Clarify-or-Direct Ans DPO model.
B.2 Additional Results
In Table 6, we report the additional results comparing different methods for determining when clarifying questions are necessary with additional base models (Llama3 and Gemma). We find that the ProCot Chain-of-Thought prompt yields degenerate behavior when used with the instruction tuned Gemma and Llama3 models, failing to ever predict to directly answer a question. When looking at the other methods, we find similar trends to the results from Llama2 discussed in Section 5.1.